OpenStudy (anonymous):

Multiply: 6 times y times the square root of 7 • 2 the square root of y. Is it 12 times y times the square root of 7 times y?

OpenStudy (anonymous):

@ivettef365?

OpenStudy (ivettef365):

so you have 6y\(\sqrt{7}\) . 2\(\sqrt{y}\)

OpenStudy (anonymous):

yup

OpenStudy (ivettef365):

and yes you are right that is 12y\(\sqrt{7y}\)

OpenStudy (anonymous):

ok thanks!

OpenStudy (ivettef365):

yw :)